caelaw:

If f(x) = 3x2 – 4 and g(x) = 2x – 6, what is g(f(2))

misty34:

There's 2 ways this can be solved. plugging the 2 into g(x) and then putting what you get from that into f(x) OR Plug g(x) into f(x) then plug in 2.

misty34:

g(x)= 2(2)-6 = 4-6= -2 f(x) = 3(-2)-4= -6-4= -10

misty34:

Your answer would be -10

misty34:

Distribute 2(2) is 4 you would put that after the = sign and -6 would go after that you take 4-6 and get -2. You plug that into f(x) Distribute 3(-2) and its -6 subtract that by -4 and you get your answer -10

misty34:

The -4 and -6 kind of tags along until you're done distributing. Make sense? I explained the easy way to do this. I hope this helps.

misty34:

@finny

finny:

change it to positive

misty34:

@jhonyy9

finny:

the negative 10 i mean

jhonyy9:

first i think may be more ussefully than you calcule f(2) = ?

jhonyy9:

please collaborate step by step

jhonyy9:

misty f(2)=?

jhonyy9:

f(2) = 3*4-4 = ?

misty34:

=8

jhonyy9:

f(2) = 8 so now just you need calcule g(8) = ? yes sure this is easy

jhonyy9:

g(8) = 2*8-6 = ?

misty34:

=10

jhonyy9:

exactly so g(f(2)) = 10

misty34:

I see. I got -10 instead but, I see what I've done wrong. Thanks for helping me fix my error

jhonyy9:

np and anytime just tag me my pleasure do you have questions ?

misty34:

I don't have any further questions

jhonyy9:

ok good luck bye bye